OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. -3, -15, -75, -375,

OpenStudy (anonymous):

an = 5 • -3^n + 1 an = -3 • 5^n - 1 an = 5 • -3^n an = -3 • 5^n

OpenStudy (anonymous):

Each successive term differs by a factor of five: -3 -3*5 -3*5*5 ... What does this look like?

OpenStudy (anonymous):

an= -3 x 5^n-1

OpenStudy (anonymous):

As long as you mean \[-3 \times5^{n-1},\text{ and not }-3\times5^n-1,\] then yes, you're right.

OpenStudy (anonymous):

ok thank you